-- FeliceCusano - 31 May 2019

INAF queue observing, 2019 May 30-31

Observer: F. Cusano
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: LBC

At the sunset the sky is clear. We start with LBC.
We wanted to change to LUCI around 8:40 UT, but for problem with instrument we stayed with LBC.

Summary

Night duration (12 degree el twilight) 3:23 - 11:11 UT = 7 h 48m

Observing time: 6h 48min (87%)

Weather time loss: 0h min (0%)

technical time loss 1h *min (13%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
calibration bias LBC-B/R       bino bias
calibration sky flat LBC-B/R g; r     good counts
Annibali R26 LBC-B/R g; r 1(bin)/1h(bin)   completed
Annibali R39 LBC-B/R g; r 1.08(bin)/1h(bin)   completed two exps more for elongation
Annibali R44 LBC-B/R g; r 0.37/0.7h(bin)   completed
Prandoni TGSSJ1530 LBC-B/R g; i 0.58/1h g; 0.58/1.5h i   last exp low elevation, but the dimm seeing was 0.5"(!)
Rossi GRB180809B LBC-B/R r; z 0.37/0.37h(bino)   completed
calibration sky flat LBC-B/R r-r; i-z     good counts

Details

01:23 starting LBC bino bias

02:30 starting sky flats g, r
  • g; r are good
  • r; i counts are too low stopped

03:06 slewing to Annibali R26, NGC4163
  • 3:13 dohybrid, too bright, we wait and chane copointing ob
  • 3:27 starting again dohybrid, no pupil found on the blue side
  • 3:33 sending again with X2
  • 3:36 dimm seeing 1"
  • 3:47 still not converging, too few pupils, the to give an offset to the telescope manually to move in a different field
  • 3:48 dofpia, /X2
  • 3:58 sending copointing, ob NGC4163_copoiting
  • 4:04 sending science, ob NGC4163_1
  • 4:03 dimm seeing 0.7"
  • 4:12 blue images look elonganted, red is ok
  • 4:18 images have not good fwhm, very high compared to dimm seeing
  • 4:20 going to a good collimation field
  • 4:26 sending again NGC4163_1
  • 4:34 images looks better average fwhm on images 0.9"
  • 4:38 right preset canceled problem with, stop the ob
  • 4:39 sending again the script starting from were it was stopped
  • 4:46 average fwhm blue side 1.0", red side 0.7"
  • 4:55 dofpia, problem with red side
  • 5:02 dohybrid
  • 5:15 NGC4163_2
  • 5:23 fwhm max on images 0.8-1"
  • 5:33 fwhm max on images 0.6-0.8"
  • 5:40 dofpia
  • 5:44 sending NGC4163_3

6:06 slewing to Annibali R39
  • 6:11 dofpia
  • 6:16 sending UGC8760_1
  • 6:16 dimm seeing 0.8"
  • 6:27 fwhm on LBC-B 1", on LBC-R 0.8"
  • 6:43 dofpia
  • 6:49 sending UGC87_60_2
  • 6:54 dimm seeing 0.7", fwhm on images ~0.7-0.8"
  • 7:18 dofpia
  • 7:21 sending UGC8760_3, two exposure on the red side are elongated, we will take two more
  • 7:50 dofpia
  • 7:5 sending UGC8760_3, for two exposure only

08:01 slewing to Annibali R44
  • 8:03 dofpia
  • 8:08 sending science NGC6789_3
  • 8:38 dimm seeing 0.9"
  • 8:39 the idea was to switch to LUCI, but there is a problem of connection with LUCI1, we stay with LBC

08:41 slewing to Prandoni, TGSSJ1530
  • 8:45 dofpia, problem with dofpia, LBC B and R not autorized but they are, we solve the problem changing account, we send dofpia from the lbto account
  • 8:58 starting science TGSSJ1530_g_i_1
  • 9:05 fwhm 0.7-1.0" on images, dimm seeing 0.7"
  • 9:20 no more space on obs2 we move the observations to obs3, no problem for the observations that keep running
  • 9:53 dofpia
  • 09:58 osa restart the software because red tracker was in a bad state, problem solved
  • 10:03 sending TGSSJ1530_g_i_2
  • 10:14 obs stopped target too low in elevation

10:14 slewing to Rossi GRB180809B
  • 10:16 dofpia
  • 10:21 starting science grb180908b_LBCbr_r+z.xml
  • 10:25 dimm seeing 0.7"
  • 10:50 dofpia
  • 10:53 sending again the script, it is still dark

11:30 starting sky flats r; i
  • r; i
  • 11:44 trying r; z flat, goo counts also *
Topic revision: r7 - 27 Sep 2019, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback